OpenStudy (anonymous):

(x^2+x-30)/(x-5) =11

OpenStudy (anonymous):

I would recommend you get rid of that fraction first of all by multiplying both sides of the equation by the denominator of the fraction (x-5) \[\frac{x^2+x-30}{x-5}\times (x-5)=11 \times (x-5)\]

OpenStudy (radar):

If you will note by inspection \[x ^{2}+x-30 =(x+6)(x-5)\]therefore the x-5 can be cancelled giving you: x+6=11

OpenStudy (anonymous):

\[\frac{x^2+x-30}{x-5} =11\]\[x^2+x-30 =11(x-5)\]\[x^2+x-30 =11x-55\]\[x^2-10x+25 =0\]\[(x-5)^2 =0\] \[x= 5\]

OpenStudy (radar):

making it simple to solve. x=5

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

Unfortunately, when x=5, the denominator x-5 is zero (since x-5=5-5=0), so x=5 cannot be a solution. So there are no solutions.

jimthompson5910 (jim_thompson5910):

This is why it's important to check any answers you get.

OpenStudy (anonymous):

Excellent catch Jim.

OpenStudy (radar):

Great spot jim, should of checked my answer.

OpenStudy (anonymous):

awesome. cause i know the answer is there is no solution because the origional equation if undefined at x=5 i just didnt know how to get there